Let ๐‘‹ be the number of times an ordinary fair die is rolled, up to and including the roll on which

the first prime number is obtained.

Find the mode of ๐‘‹

Find ๐ธ(๐‘‹)

Evaluate ๐‘ƒ(๐‘‹ > ๐ธ(๐‘‹))


1
Expert's answer
2021-01-11T18:10:50-0500

"prime\;numbers\;that\; we\;can\;obtain\;are\;2,3,5\\
\therefore p=0.5,\\
P(x)=0.5(0.5)^{x-1},x=0,1,...\\
Mode=1,\\
E(x)=1/p=1/0.5=2,\\
𝑃(𝑋 > 𝐸(𝑋))=𝑃(𝑋 > 2)\\
=1-P(X\leq 2)=1-[P(X=1)+P(X=2)]\\
=1-[0.5(0.5)^{1-1}+0.5(0.5)^{2-1}]=0.25"